OpenStudy (anonymous):

the ratio of8:5 = (?)% please break it down to the me in the easiest terms

hero (hero):

8:5 = 8/5 =

OpenStudy (anonymous):

3/5 = (?)% * 5/8

OpenStudy (anonymous):

8/5 =8:5=1.6=160%

OpenStudy (angela210793):

8/5*100=160%